Calculate the pH at each point listen for the titration of 50.0 mL of 0.050 0 M formic acid with 0.050 0 M KOH. The points to calculate are Vb = 0.0, 10.0, 20.0, 25.0, 30.0, 40.0, 45.0, 48.0, 49.0, 49.5, 50.0, 50.5, 51.0, 52.0, 55.0, and 60.0 mL. Draw a graph of pH versus Vb.


1
Expert's answer
2021-10-07T01:16:01-0400

pH= pKa+log(acid/base)

Molar Mass of KOH = 56.1056 g/mol

56.1056×0.05= 2.81g

50×0.05 =0.25

Log(2.81/0.25)

= 1.05+pKa

1.05+3.74 = 4.79